# Hyperbolic Graph Learning: A Comprehensive Review

### v3

@@ 2.1 Mathematical Preliminaries
[…]
For each point $\mathbf{x}$ in $\mathcal{M}$, a tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}$ is defined as the first-order approximation of $\mathcal{M}$ around $\mathbf{x}$, which is an $n$-dimension vector space and isomorphic to $\mathbb{R}^{n}$. The Riemannian manifold metric $g^{\mathcal{M}}$ assigns a smoothly varying positive definite inner product $<\cdot,\cdot>:\mathcal{T}_{\mathbf{x}}\mathcal{M}\times\mathcal{T}_{\mathbf{x}}\mathcal{M}\to\mathbb{R}$ on the tangent space, which allows us to define several geometric properties, such as geodesic distances, angles, and curvature.
[…]
@@ 2.2 Isometric Models in the Hyperbolic Space
[…]
@@ 3 Hyperbolic Graph Representation Methods
   Mechanism (§3.3.3), middle, text width=12.0em, edge=output-black [ HGNN [37], $\kappa$HGCN [62], HGCN [36], HAT [38], LGCN [57],
@@ 3.2.2 Angle-based Methods
[…]
@@
$\left.\mathfrak{S}_{x}^{\psi(x)}=\left\{\begin{array}[]{cc}y\in\mathcal{B}^{n}&\Xi(x,y)\leq\arcsin\left(K\frac{1-\|x\|^{2}}{\|x\|}\right)\end{array}\right.\right\},$ (13)
@@
$\overline{\mathbf{h}}_{i}=\widehat{\mathbf{h}}_{i}/||\widehat{\mathbf{h}}_{i}||,\widehat{\mathbf{h}}_{i}=\log_{\mathbf{h}_{i}}(\frac{1+||\mathbf{h}_{i}||}{2||\mathbf{h}_{i}||}\mathbf{h}_{i}),$ (19)
@@ 3.3.2 Fully Hyperbolic Models
$\mathbf{W}{\otimes^{\mathcal{B}}}\mathbf{x}^{\mathcal{B}}=\tanh\left(\frac{||\mathbf{W}\mathbf{x}^{\mathcal{B}}||}{||\mathbf{x}^{\mathcal{B}}||}\mathrm{actanh}(||\mathbf{x}^{\mathcal{B}}||)\right)\frac{\mathbf{W}\mathbf{x}^{\mathcal{B}}}{||\mathbf{W}\mathbf{x}^{\mathcal{B}}||},$ (28)
@@
$\mathbf{x}^{\mathcal{B}}\oplus^{\mathcal{B}}\mathbf{y}^{\mathcal{B}}=\frac{(1+2\langle\mathbf{x}^{\mathcal{B}},\mathbf{y}^{\mathcal{B}}\rangle+||\mathbf{y}^{\mathcal{B}}||^{2})\mathbf{x}^{\mathcal{B}}+(1-||\mathbf{x}^{\mathcal{B}}||^{2})\mathbf{y}^{\mathcal{B}}}{1+2\langle\mathbf{x}^{\mathcal{B}},\mathbf{y}^{\mathcal{B}}\rangle+||\mathbf{x}^{\mathcal{B}}||^{2}||\mathbf{y}^{\mathcal{B}}||^{2}}.$ (29)
@@
where $\mathbf{\alpha}=(\alpha_{1},...,\alpha_{N})$ as the weights for each sample $\mathbf{x}^{\mathcal{B}}_{i}$ and $\gamma_{i}=\frac{2}{||\mathbf{x}^{\mathcal{B}}_{i}||^{2}}$. These three centroids can be characterized as a minimizer of the weighted sum of calibrated squared distance [78].
@@ 3.3.3 Neighborhood Aggregation Mechanism
The determination of neighborhood weights is crucial in aggregating information effectively. Various strategies can be employed to calculate these weights, including utilizing structure information, feature information, hyperbolic distance, or a combination of these methods, which are summarized in Table II. Below, we delve into each of these strategies. Consider a central node $i$, and we assume that ${N}(i)$ represents the immediate one-hop neighbor of node $i$, with $j$ being an element of $N(i)$ [^3] Following the literature, we consider $i\in{N}(i)$..

### v4

# Hyperbolic Graph Neural Networks: A Review of Methods and Applications

@@ 2.1 Mathematical Preliminaries
Note. The Klein norm at point $\mathbf{x}$ is given by: $\|\mathbf{v}\|_{\mathbf{x}}^{\mathcal{K}}=\sqrt{\frac{\|\mathbf{v}\|^{2}-c(\langle\mathbf{x},\mathbf{v}\rangle)^{2}}{1+c\|\mathbf{x}\|^{2}}}$; The Klein inner product at point $\mathbf{x}$ is given by: $\langle\mathbf{u},\mathbf{v}\rangle_{\mathbf{x}}^{\mathcal{K}}=\frac{\langle\mathbf{u},\mathbf{v}\rangle-c\langle\mathbf{x},\mathbf{u}\rangle\langle\mathbf{x},\mathbf{v}\rangle}{1+c\|\mathbf{x}\|^{2}}$; All inner products $\langle\cdot,\cdot\rangle$ without a superscript refer to standard Euclidean inner products.
For each point $\mathbf{x}$ in $\mathcal{M}$, a tangent space $\mathcal{T}_{\mathbf{x}}\mathcal{M}$ is defined as the first-order approximation of $\mathcal{M}$ around $\mathbf{x}$, which is an $n$-dimension vector space and isomorphic to $\mathbb{R}^{n}$. The Riemannian manifold metric $g^{\mathcal{M}}$ assigns a smoothly varying positive definite inner product $\cdot,\cdot>:\mathcal{T}_{\mathbf{x}}\mathcal{M}\times\mathcal{T}_{\mathbf{x}}\mathcal{M}\to\mathbb{R}$ on the tangent space, which allows us to define several geometric properties, such as geodesic distances, angles, and curvature.
Geodesics and Induced Distance Function. For a curve $\gamma:[\alpha,\beta]\to\mathcal{M}$, the shortest length of $\gamma$, i.e., geodesics, is defined as $L(\gamma)=\int_{\alpha}^{\beta}\|\gamma^{\prime}(t)\|_{g}dt$. Then the distance of $\mathbf{u},\mathbf{v}\in\mathcal{M}$, $d_{\mathcal{M}}(\mathbf{u},\mathbf{v})=\inf L(\gamma)$ where $\gamma$ is a curve that $\gamma(\alpha)=\mathbf{u},\gamma(\beta)=\mathbf{v}$.
@@ 2.2 Isometric Models in the Hyperbolic Space
With negative curvature $c\quad(c<0)$, the Lorentz model (also named hyperboloid model) defines $n$-dimensional hyperbolic space as the upper sheet of a two-sheeted hyperboloid embedded in $(n+1)$-dimensional Minkowski space. Minkowski space is equipped with the Minkowski inner product $\left\langle\mathbf{x},\mathbf{y}\right\rangle_{\mathcal{L}}=-x_{0}y_{0}+\sum_{i=1}^{n}x_{i}y_{i}$. The hyperboloid is defined as the Riemannian manifold $(\mathcal{L}_{c}^{n},g_{\mathbf{x}}^{\mathcal{L}_{c}^{n}})$, where $\mathcal{L}_{c}^{n}=\left\{\mathbf{x}\in\mathbb{R}^{n+1}:\left\langle\mathbf{x},\mathbf{x}\right\rangle_{\mathcal{L}}={1}/{c}\right\}$ and its metric tensor $g_{\mathbf{x}}^{\mathcal{L}_{c}^{n}}=\frac{1}{|c|}\cdot\text{diag}([-1,1,...,1])_{n}$.
@@ 3 Hyperbolic Graph Representation Methods
[…]
@@ 3.2.2 Angle-based Methods
To avoid heavy cone intersections and ensure capacity that scales exponentially with the dimensionality of the space, angular cones are constructed within the Poincaré ball. The angular cones meet four intuitive properties, including axial symmetry, rotation invariance, continuity of cone aperture function, and the transitivity of nested angular cones. HECone [51] further provided a closed-form of the angular cones, for any $\mathbf{x},\mathbf{y}\in\mathcal{B}^{n}\setminus\mathcal{B}^{n}(O,\epsilon)$, the angle between the half-lines (xy({x}{y} and (Ox({O}{x} as:
@@
[…]
@@
[…]
@@ 3.3.2 Fully Hyperbolic Models
[…]
@@
[…]
@@
[…]
@@ 3.3.3 Neighborhood Aggregation Mechanism
The determination of neighborhood weights is crucial in aggregating information effectively. Various strategies can be employed to calculate these weights, including utilizing structure information, feature information, hyperbolic distance, or a combination of these methods, which are summarized in Table II. Below, we delve into each of these strategies. Consider a central node $i$, and we assume that N(i){N}(i) represents the immediate one-hop neighbor of node $i$, with $j$ being an element of N(i)N(i) [^3] Following the literature, we consider $i\in{N}(i)$..